# Spatial Blind Spot: Auditory Motion Perception Deficits in Audio LLMs

### v1

@@
Abstract
Large Audio-Language Models (LALMs) have recently shown impressive progress in speech recognition, audio captioning, and auditory question answering. Yet, whether these models can perceive spatial dynamics, particularly the motion of sound sources, remains unclear. In this work, we uncover a systematic motion perception deficit in current ALLMs. To investigate this issue, we introduce AMPBench, the first benchmark explicitly designed to evaluate auditory motion understanding. AMPBench introduces a controlled question-answering benchmark designed to evaluate whether Audio-Language Models (LALMs) can infer the direction and trajectory of moving sound sources from binaural audio. Comprehensive quantitative and qualitative analyses reveal that current models struggle to reliably recognize motion cues or distinguish directional patterns. The average accuracy remains below 50%, underscoring a fundamental limitation in auditory spatial reasoning. Our study highlights a fundamental gap between human and model auditory spatial reasoning, providing both a diagnostic tool and new insight for enhancing spatial cognition in future Audio-Language Models.
@@ 1 Introduction
Recent Large Audio-Language Models (LALMs), such as Qwen2-Audio-7B, Aero-1-Audio, Voxtral-Mini-3B-2507,and Kimi-Audio-7B [chu2024qwen2, liu2025voxtral, mandel2023aero, ding2025kimi], have achieved remarkable progress in understanding and reasoning over speech, music, and environmental sounds. Yet a fundamental aspect of auditory intelligence remains largely unexplored: the perception of motion. When these models are asked a seemingly simple question, “Which way is the sound moving?”, they hesitate, guess, or fail. Figure 1 presents a concrete case that highlights this difficulty. In the example, a tiger moves around a small robot representing an audio–language model. The robot can readily identify that the audio contains a tiger vocalizing, yet it cannot describe the direction or trajectory of the tiger’s movement. The accompanying binaural waveforms further illustrate the richness of spatial information that current LALMs fail to exploit. Each of the four panels corresponds to a distinct motion trajectory of the tiger, and the left–right amplitude patterns encode clear interaural cues. For example, in the left-to-right motion, the left channel exhibits an early high-energy onset followed by a gradual transition toward the right channel, whereas in trajectories such as left-front to right-front or right-front to right, the amplitude envelope and decay pattern reflect coordinated changes in both lateral position and distance as the tiger moves. However, despite the abundance of physically meaningful spatial cues embedded in these signals, the model is unable to convert them into a coherent estimate of the tiger’s direction or path.
Figure 1: Illustration of the motion perception gap in LALMs. The upper-left panel depicts a tiger moving along four distinct trajectories around the model, and the right panels show the corresponding binaural waveforms. Each trajectory yields characteristic left–right patterns, including asymmetric amplitude envelopes, interaural intensity differences, and distance-dependent decay that jointly encode the underlying spatial motion. Despite the richness of these spatial cues, current LALMs fail to infer its direction or trajectory.
@@
Human auditory intelligence, by contrast, is profoundly spatial. The brain continuously fuses interaural time differences (ITD), interaural level differences (ILD), Doppler shifts, and spectral filtering to reconstruct the three-dimensional motion of sound sources [chu2025learning]. Such spatial integration enables humans to localize, track, and anticipate motion even in noisy or reverberant conditions. In contrast, most existing LALMs are trained on single-channel or weakly spatialized datasets that emphasize semantic comprehension but neglect spatial reasoning. As a result, their learned representations capture semantics without geometry, leading to a fundamental question: Can modern Audio-Language Models truly perceive motion, or are they merely predicting words about sounds?
@@
To answer this question, we introduce the Audio Motion Perception Benchmark (AMPBench), the first benchmark specifically designed to evaluate motion understanding in LALMs. AMPBench provides physically simulated binaural audio clips under controlled spatial conditions, paired with structured question-answering tasks that probe directionality and trajectory reasoning across multiple noise levels. This benchmark moves beyond static sound recognition to test whether models can infer spatial dynamics directly from the waveform. Through extensive zero-shot evaluation of five state-of-the-art systems, we reveal a consistent and systematic motion perception deficit: models fail to correctly recognize motion cues and directional patterns, achieving below 50% accuracy on average. This benchmark thus exposes a key limitation in current audio-language understanding and establishes a diagnostic foundation for developing spatially aware auditory models.
@@
To address these limitations, we obtained the following research results through controlled experiments and systematic benchmark evaluations:
@@
  • We propose AMPBench, a controlled and reproducible benchmark for evaluating spatial reasoning from binaural audio. It provides binaural simulations under well-defined spatial parameters and structured question–answering tasks that directly test models’ understanding of auditory motion.
@@ 2.1 Large Audio-Language Models
Large Audio-Language Models (LALMs) extend large language models by incorporating audio perception and reasoning capabilities. These models typically employ pretrained audio encoders (e.g., WavLM[chen2022wavlm]) to extract high-level acoustic representations, which are then fused with a language-model backbone and fine-tuned for downstream audio-text tasks [su2025audio]. Recent surveys show the field of LALMs is growing rapidly, covering model architecture, multimodal fusion strategies, and evaluation benchmarks [yang2025towards].
Several recent systems exemplify this trend[ding2025kimi, xie2025audio, hurst2024gpt, ghosh2024gama, goel2025audio]. For example, the Qwen-Audio model unified pre-training across a large set of audio tasks and domains and achieved strong zero-shot performance on audio understanding benchmarks[chu2023qwen]. More recently, the Qwen2.5-Omni model pushes toward full multimodal input including audio, video and image, thus marking a transition from pure audio-language to richer modality integration [xu2025qwen2]. Meanwhile, Xiaomi’s MiMo-Audio model—trained on over 100 million hours of audio and using a unified audio-text next-token objective—demonstrates the push for scale and generality in audio-language modelling [coreteam2025mimoaudio].
@@ 2.2 Spatial and Multimodal Perception
Spatial and multimodal perception in the audio domain addresses the “where” and “how” of sound: not only what source is present, but also its direction, distance, motion, and relationship with visual/contextual cues. Recent research has begun to explore spatial audio understanding and generation, highlighting the need for models that go beyond semantic recognition and reason about spatial and multimodal context [zhu2025asaudio, sarabia2023spatial].
For example, the ELSA model (Embeddings for Language and Spatial Audio) learns spatially-aware audio and text embeddings via contrastive learning over spatial configurations, showing that standard audio‐text models lack spatial awareness [devnani2024learning]. The BAT model (Learning to Reason about Spatial Sounds with Large Language Models) demonstrates how a spatial audio encoder can be combined with a language model to answer spatial sound reasoning questions [zheng2024bat]. More recently, SALM (Spatial Audio Language Model) integrates structured audio embeddings separating semantic and spatial components, enabling zero-shot direction classification and editing of spatial audio scenes [hu2025salm]. These works illustrate a growing trend toward spatially-grounded audio-language modelling.
Meanwhile, Sarabia et al. introduced Spatial LibriSpeech, a dataset augmented with multi-channel spatial audio and annotations for source position, direction and room acoustics [sarabia2023spatial]. Other works examine spatially aligned audio–video generation benchmarks such as SAVGBench which focus on spatialized audio–visual coherence [shimada2024savgbench].
Despite this progress, most LALMs still rely on single-channel or weakly spatialized datasets and rarely model motion trajectories or multimodal cues (e.g., visual context) in their training. The field of spatial audio generation further shows that tasks such as moving sound source synthesis, multichannel rendering, and binaural reasoning remain under-explored [zhu2025asaudio, liu2025text2move, groger2025evaluation]. Consequently, spatial and multimodal perception remains a bottleneck for truly embodied audio-language intelligence.
@@ 2.3 Benchmarks and Datasets
Several benchmarks have been developed to evaluate audio–language models, yet few are tailored explicitly for spatial or motion reasoning in auditory scenes. For instance, the AudioBench benchmark provides a universal evaluation framework covering eight tasks and 26 datasets targeting LALMs [wang2024audiobench]. The MMAU benchmark further advances this by offering 10 000 audio clips paired with human-annotated questions and answers spanning speech, environmental sounds and music, and evaluates models on 27 distinct skills including reasoning and information extraction [sakshi2024mmau]. Another recent benchmark, AIR-Bench, is designed to assess LALMs’ interaction capabilities via generative comprehension across speech, natural sounds and music, with both foundation and chat tasks [yang2024air].
While these resources provide valuable coverage of semantic and reasoning tasks, they largely overlook the continuous spatial and motion aspects of auditory scenes. This gap motivates our proposed benchmark designed specifically for auditory motion perception and reasoning from binaural audio.
3 The AMPBench Framework
Building upon the limitations identified in prior benchmarks, AMPBench aims to provide a systematic evaluation of auditory motion perception. Whereas previous datasets emphasize semantic or multimodal understanding, AMPBench isolates motion as an independent perceptual dimension and measures whether large audio–language models (LALMs) can infer direction and continuity purely from binaural cues. The following subsections describe its task formulation, synthesis pipeline, and dataset composition in detail.
@@ 3.1 Task Formulation
AMPBench evaluates a model’s ability to perceive and reason about moving sound sources using only auditory input. We formulate this as an auditory motion question–answering task in which a model must infer the direction and trajectory of a moving sound source based on a binaural waveform and a natural-language query.
Existing audio benchmarks occasionally contain motion-related traces, yet such cues are typically incidental and irregular. The temporal evolution of interaural cues such as ITD and ILD is neither controlled nor uniformly covered across directions, and motion trajectories are seldom parameterized for reproducibility. Moreover, real-world audio is often acoustically complex: changes in timbre, overlapping events, and reverberation introduce non-spatial information that can mislead a model’s judgment. As a result, previous datasets primarily reflect semantic recognition or coarse spatial tagging, but rarely examine whether a model can truly understand dynamic soundfields through continuous auditory evidence rather than linguistic or incidental acoustic cues.
AMPBench closes this gap by isolating spatial–temporal reasoning from semantic understanding. It defines a minimal label-constrained question–answering task in which binaural waveforms are synthesized under physically controlled parameters, ensuring that spatial cues dominate while non-spatial confounds are minimized. By enforcing balanced coverage over canonical start–end positions, explicitly controlling pitch and speed, and standardizing prompts to remove linguistic bias, AMPBench provides a physically grounded and evaluation-stable benchmark for testing whether LALMs can infer motion direction and continuity directly from binaural perception.
@@
$\begin{gathered}f_{\theta}(x,q)\rightarrow y,\\ y\in\begin{cases}\mathcal{Y}_{\mathrm{MC}}=\{\mathrm{A,B,C,D}\},&\text{multiple choice},\\[2.0pt] \mathcal{Y}_{\mathrm{TF}}=\{\mathrm{TRUE,FALSE}\},&\text{true/false}.\end{cases}\end{gathered}$ (1)
@@ 3.3 Benchmark Composition and Variants
The AMPBench dataset systematically covers all combinations of the eight canonical spatial positions, resulting in a total of $8\times 7=56$ directed motion trajectories. Each trajectory corresponds to a physically simulated sound source traveling linearly from one position to another within the $6\,\text{m}\times 6\,\text{m}$ spatial field, with the listener fixed at the center $(3.0,3.0)$. With an average clip duration of approximately $6\,\text{s}$ determined by the rendering pipeline, the full set of 56 trajectories yields a total of about $5.6$ minutes of spatial audio. This exhaustive pairing ensures balanced sampling across lateral, diagonal, and front–back motions, providing comprehensive coverage of the auditory motion space.
For each trajectory, binaural recordings are generated under four acoustic conditions: one clean (noise-free) and three levels of additive Gaussian noise corresponding to approximate signal-to-noise ratios of 35, 25, and 15 dB. Thus, every directional path yields four distinct versions of the same motion sequence, facilitating controlled robustness evaluation. In total, AMPBench contains $56\times 4=224$ synthesized audio clips.
Figure 2: Examples from AMPBench, including (a) a multiple-choice question and (b) a true–false verification. Each item couples a synthesized binaural motion clip with a structured QA prompt for label-based evaluation.
@@
To enable fine-grained analysis, AMPBench further defines three controlled task variants probing complementary reasoning dimensions. The first variant, Fixed Pitch Motion, keeps both frequency and velocity constant to isolate ITD/ILD-based spatial perception. The second, Variable Pitch Motion, varies the source pitch among seven discrete frequencies to test the ability to disentangle spatial cues from spectral interference. The third, Variable Speed Motion, maintains a fixed pitch while adjusting the motion velocity by a controllable scaling factor (e.g., doubled speed) to examine temporal adaptability and the continuity of motion reasoning. Together, these variants jointly assess the spatial, spectral, and temporal aspects of auditory reasoning under matched experimental protocols.
Table 1: Overview of the three AMPBench task variants, each targeting a distinct perceptual dimension of auditory motion.
@@
Fixed Pitch Constant pitch and velocity across trajectories. 224 MCQ + 448 T/F
Variable Pitch Pitch varies among seven discrete frequencies. 224 MCQ + 448 T/F
Variable Speed Fixed pitch but adjustable velocity. 224 MCQ + 448 T/F
By combining physical simulation, linguistic structure, and controlled variability, AMPBench offers a comprehensive and interpretable testbed for auditory motion reasoning. It bridges low-level binaural cues with high-level question-answering evaluation—forming a foundation for the next generation of spatially aware audio–language models.

### v2

# AudioMotionBench: Evaluating Auditory Motion Perception in Audio LLMs

@@
Abstract Large Audio-Language Models (LALMs) have recently shown impressive progress in speech recognition, audio captioning, and auditory question answering. Yet, whether these models can perceive spatial dynamics, particularly the motion of sound sources, remains unclear. In this work, we uncover a systematic motion perception deficit in current ALLMs. To investigate this issue, we introduce AudioMotionBench, the first benchmark explicitly designed to evaluate auditory motion understanding. AudioMotionBench introduces a controlled question-answering benchmark designed to evaluate whether Audio-Language Models (LALMs) can infer the direction and trajectory of moving sound sources from binaural audio. Comprehensive quantitative and qualitative analyses reveal that current models struggle to reliably recognize motion cues or distinguish directional patterns. The average accuracy remains below 50%, underscoring a fundamental limitation in auditory spatial reasoning. Our study highlights a fundamental gap between human and model auditory spatial reasoning, providing both a diagnostic tool and new insight for enhancing spatial cognition in future Audio-Language Models.
@@ 1 Introduction
Figure 1: Overview of AudioMotionBench and the audio motion perception evaluation pipeline,The benchmark simulates controlled binaural audio with explicit motion cues to model moving sound sources in realistic environments. Dual-channel audio inputs are encoded and paired with structured multiple-choice questions that require reasoning about sound source direction and trajectory.
@@
Recent Large Audio-Language Models (LALMs), such as Qwen2-Audio-7B, Aero-1-Audio, Voxtral-Mini-3B-2507,and Kimi-Audio-7B chu2024qwen2, liu2025voxtral, mandel2023aero, ding2025kimi, have achieved remarkable progress in understanding and reasoning over speech, music, and environmental sounds. Yet a fundamental aspect of auditory intelligence remains largely unexplored: the perception of motion. While current models perform well on recognizing sound types and semantic content, existing audio-language tasks rarely require them to reason about the dynamic movement and trajectory of sound sources in space.
@@
Human auditory intelligence, is profoundly spatial. The brain continuously fuses interaural time differences (ITD), interaural level differences (ILD), Doppler shifts, and spectral filtering to reconstruct the three-dimensional motion of sound sources chu2025learning. Such spatial integration enables humans to localize, track, and anticipate motion even in noisy or reverberant conditions. In contrast, most existing LALMs are trained on single-channel or weakly spatialized datasets that emphasize semantic comprehension but neglect spatial reasoning, leading to a fundamental question: Can modern Audio-Language Models truly perceive motion?
@@
To answer this question, we introduce the Audio Motion Perception Benchmark (AudioMotionBench) (Figure 1), the first benchmark specifically designed to evaluate motion understanding in LALMs. AudioMotionBench provides physically simulated binaural audio clips under controlled spatial conditions, paired with structured question-answering tasks that probe directionality and trajectory reasoning across multiple noise levels. This benchmark moves beyond static sound recognition to test whether models can infer spatial dynamics directly from the waveform. Through extensive zero-shot evaluation of five state-of-the-art systems, we reveal a consistent and systematic motion perception deficit: models fail to correctly recognize motion cues and directional patterns, achieving below 50% accuracy on average. This benchmark thus exposes a key limitation in current audio-language understanding and establishes a diagnostic foundation for developing spatially aware auditory models.
@@
We obtained the following research results through controlled experiments and systematic benchmark evaluations:
@@
  • We propose AudioMotionBench, a controlled and reproducible benchmark for evaluating spatial reasoning from binaural audio. It provides binaural simulations under well-defined spatial parameters and structured question–answering tasks that directly test models’ understanding of auditory motion.
@@ 2.1 Large Audio-Language Models
Large Audio-Language Models (LALMs) extend large language models by incorporating audio perception and reasoning capabilities. These models typically employ pretrained audio encoders (e.g., WavLMchen2022wavlm) to extract high-level acoustic representations, which are then fused with a language-model backbone and fine-tuned for downstream audio-text tasks su2025audio. Recent surveys show the field of LALMs is growing rapidly, covering model architecture, multimodal fusion strategies, and evaluation benchmarks yang2025towards.
Several recent systems exemplify this trendding2025kimi, xie2025audio, hurst2024gpt, ghosh2024gama, goel2025audio. For example, the Qwen-Audio model unified pre-training across a large set of audio tasks and domains and achieved strong zero-shot performance on audio understanding benchmarkschu2023qwen. More recently, the Qwen2.5-Omni model pushes toward full multimodal input including audio, video and image, thus marking a transition from pure audio-language to richer modality integration xu2025qwen2. Meanwhile, Xiaomi’s MiMo-Audio model—trained on over 100 million hours of audio and using a unified audio-text next-token objective—demonstrates the push for scale and generality in audio-language modelling coreteam2025mimoaudio.
@@ 2.2 Spatial and Multimodal Perception
Spatial and multimodal perception in the audio domain addresses the “where” and “how” of sound: not only what source is present, but also its direction, distance, motion, and relationship with visual/contextual cues. Recent research has begun to explore spatial audio understanding and generation, highlighting the need for models that go beyond semantic recognition and reason about spatial and multimodal context zhu2025asaudio, sarabia2023spatial.
For example, the ELSA model (Embeddings for Language and Spatial Audio) learns spatially-aware audio and text embeddings via contrastive learning over spatial configurations, showing that standard audio‐text models lack spatial awareness devnani2024learning. The BAT model (Learning to Reason about Spatial Sounds with Large Language Models) demonstrates how a spatial audio encoder can be combined with a language model to answer spatial sound reasoning questions zheng2024bat. More recently, SALM (Spatial Audio Language Model) integrates structured audio embeddings separating semantic and spatial components, enabling zero-shot direction classification and editing of spatial audio scenes hu2025salm. These works illustrate a growing trend toward spatially-grounded audio-language modelling.
Meanwhile, Sarabia et al. introduced Spatial LibriSpeech, a dataset augmented with multi-channel spatial audio and annotations for source position, direction and room acoustics sarabia2023spatial. Other works examine spatially aligned audio–video generation benchmarks such as SAVGBench which focus on spatialized audio–visual coherence shimada2024savgbench.
Despite this progress, most LALMs still rely on single-channel or weakly spatialized datasets and rarely model motion trajectories or multimodal cues (e.g., visual context) in their training. The field of spatial audio generation further shows that tasks such as moving sound source synthesis, multichannel rendering, and binaural reasoning remain under-explored zhu2025asaudio, liu2025text2move, groger2025evaluation. Consequently, spatial and multimodal perception remains a bottleneck for truly embodied audio-language intelligence.
@@ 2.3 Benchmarks and Datasets
Several benchmarks have been developed to evaluate audio–language models. For instance, the AudioBench benchmark provides a universal evaluation framework covering eight tasks and 26 datasets targeting LALMs wang2024audiobench. The MMAU benchmark further advances this by offering 10,000 audio clips paired with human-annotated questions and answers spanning speech, environmental sounds and music, and evaluates models on 27 distinct skills including reasoning and information extraction sakshi2024mmau. Another recent benchmark, AIR-Bench, is designed to assess LALMs’ interaction capabilities via generative comprehension across speech, natural sounds and music, with both foundation and chat tasks yang2024air.
While these resources provide valuable coverage of semantic and reasoning tasks, they largely overlook the continuous spatial and motion aspects of auditory scenes.As summarized in Table 1, existing audio–language benchmarks primarily emphasize semantic understanding and broad task coverage, while largely overlooking controlled auditory motion perception. In contrast, AudioMotionBench is explicitly designed to isolate continuous spatial cues in binaural audio under physically controlled conditions, minimizing semantic and acoustic confounds. This makes AudioMotionBench complementary to prior benchmarks and uniquely suited for probing motion-specific auditory reasoning.
Benchmark Audio Modality Task Diversity Continuous Spatial Cues Real-World Audio
AudioBench wang2024audiobench Mono / Multi High
MMAU sakshi2024mmau Mono High
AIR-Bench yang2024air Mono Medium
AudioMotionBench Binaural Medium
Table 1: Comparison between AudioMotionBench and existing audio–language benchmarks. Existing benchmarks emphasize broad task coverage and real-world audio, while AudioMotionBench focuses on controlled auditory motion perception through continuous spatial cues in binaural signals.
3 The AudioMotionBench Framework
Building upon the limitations identified in prior benchmarks, AudioMotionBench aims to provide a systematic evaluation of auditory motion perception. Whereas previous datasets emphasize semantic or multimodal understanding, AudioMotionBench isolates motion as an independent perceptual dimension and measures whether large audio–language models (LALMs) can infer direction and continuity purely from binaural cues. The following subsections describe its task formulation, synthesis pipeline, and dataset composition in detail.
@@ 3.1 Task Formulation
AudioMotionBench evaluates a model’s ability to perceive and reason about moving sound sources using only auditory input. We formulate this as an auditory motion question–answering task in which a model must infer the direction and trajectory of a moving sound source based on a binaural waveform and a natural-language query.
@@
$\begin{gathered}y=f_{\theta}(x,q),\\ y\in\begin{cases}\mathcal{Y}_{\mathrm{MC}}=\{\mathrm{A,B,C,D}\},&\text{multiple choice},\\[2.0pt] \mathcal{Y}_{\mathrm{TF}}=\{\mathrm{TRUE,FALSE}\},&\text{true/false}.\end{cases}\end{gathered}$ (1)
@@ 3.3 Benchmark Composition and Variants
The AudioMotionBench systematically covers all combinations of the eight canonical spatial positions, resulting in a total of $8\times 7=56$ directed motion trajectories. Each trajectory corresponds to a physically simulated sound source traveling linearly from one position to another within the $6\,\text{m}\times 6\,\text{m}$ spatial field, with the listener fixed at the center $(3.0,3.0)$. With an average clip duration of approximately $6\,\text{s}$ determined by the rendering pipeline, the full set of 56 trajectories yields a total of about $5.6$ minutes of spatial audio. This exhaustive pairing ensures balanced sampling across lateral, diagonal, and front–back motions, providing comprehensive coverage of the auditory motion space.
For each trajectory, binaural recordings are generated under four acoustic conditions: one clean (noise-free) and three levels of additive Gaussian noise corresponding to approximate signal-to-noise ratios of 35, 25, and 15 dB. Thus, every directional path yields four distinct versions of the same motion sequence, facilitating controlled robustness evaluation. In total, AudioMotionBench contains $56\times 4=224$ synthesized audio clips.
Multiple-choice Question Prompt:
Which option best describes the motion direction?
Answer with one letter (A/B/C/D).
A  From front-left to back-left B  From back-left to front-left C  From front-right to front-left D  From front-left to front-right  ✓
True/False Verification Prompt:
Listen to the following binaural audio and judge whether the statement is correct.
Answer only with “True” or “False”.
TRUE  The statement is correct. FALSE  The statement is wrong.  ✓
Figure 2: Example question formats in AudioMotionBench, including well-designed multiple-choice and true/false questions.
@@
To enable fine-grained analysis, AudioMotionBench further defines three controlled task variants probing complementary reasoning dimensions. The first variant, Spatial Consistency Task (SCT), keeps both frequency and velocity constant to isolate ITD/ILD-based spatial perception. The second, Spectral Disentanglement Task (SDT), varies the source pitch among seven discrete frequencies to test the ability to disentangle spatial cues from spectral interference. The third, Temporal Adaptation Task (TAT), maintains a fixed pitch while adjusting the motion velocity by a controllable scaling factor (e.g., doubled speed) to examine temporal adaptability and the continuity of motion reasoning. Together, these variants jointly assess the spatial, spectral, and temporal aspects of auditory reasoning under matched experimental protocols.
@@
SCT Constant pitch and velocity. 224 MCQ + 448 T/F
SDT Discrete pitch variation. 224 MCQ + 448 T/F
TAT Fixed pitch, variable velocity. 224 MCQ + 448 T/F
Table 2: Overview of AudioMotionBench task variants.
By combining physical simulation, linguistic structure, and controlled variability, AudioMotionBench offers a comprehensive and interpretable testbed for auditory motion reasoning. It bridges low-level binaural cues with high-level question-answering evaluation—forming a foundation for the next generation of spatially aware audio–language models.